With the gasoline time series data from the given table, show the exponential smoothing forecasts using a = 0.1. GASOLINE SALThe options for prefer for A B and C are 0.1 or 0.2
Answer #1

For Alpha=0.1 the calculations are shown as

17 17 17 - 17 = 0 (17- 17) = 0 17-11= 0% 21-17 21 17 21 - 17 = 4 19.05% (21 - 17) = 16 (19- 17.4) = 2.56 19 17.4 19 - 17.4 =

For alpha = 0.2 the calculation is shown as

Data ES(0.2) Forecasts Abs. Error Error2 Abs. % Error 17 17 17 - 17 = 0 12 = 0% (17- 17) = 0 21 - 17) = 16 21 - 17 = 21 17 4

MAE a=0.1 MSE 8.48 MAPE 11.87 2.35 æ0.2 2.38 8.23 2-29 acc tee MS @=0.2 ace tee MAE a=0.1 pro. I acc ta MAPE